OpenStudy (anonymous):

Divide x to the 1 half power divided by x to the 2 sevenths power I'm stuck on this one! plz help. Its for Algebra 2

OpenStudy (johnweldon1993):

\[\huge \frac{x^{\frac{1}{2}}}{x^{\frac{2}{7}}}\]

OpenStudy (anonymous):

[\sqrt{x}/\sqrt[7]{x ^{2}}\]

OpenStudy (johnweldon1993):

When dividing exponents with the same base...you leave the base (x) alone...and just subtract the exponents \[\huge x^{(\frac{1}{2}-\frac{2}{7}})\]

OpenStudy (anonymous):

\[\sqrt{x}\div \sqrt[7]{x ^{2}}\]

OpenStudy (anonymous):

remember, when x is in the denominator, it can easily be brought up to the numerator by changing the sign of it's exponent. \[x ^{-\frac{ 2 }{ 7 }}\] in the numerator and take it from there by using the "same base" rule

OpenStudy (anonymous):

\[\frac{ x ^{\frac{ 1 }{2}} }{ x ^{\frac{ 2 }{ 7 }} }=x ^{\frac{ 1 }{ 2 }}*x ^{\frac{ -2 }{ 7 } } =x ^{\left( \frac{1 }{ 2 }-\frac{ 2 }{7 } \right)}=x ^{\frac{ 1*7-2*2 }{ 2*7 }}=x ^{\frac{ 3 }{ 14 }}\]